Bobo can wash the car in 36 minutes by himself. When his brother Otto helps, they finish in 20 minutes. How long would it take Otto to wash the car by himself?
Answer by stanbon(75887)

Bobo can wash the car in 36 minutes by himself. When his brother Otto helps, they finish in 20 minutes. How long would it take Otto to wash the car by himself?
----
Bobo rate = 1/36 job/min
----
Together rate = 1/20 job/min
---
Otto rate = 1/x job/min
---------------------------
Equation:
rate + rate = together rate
1/36 + 1/x = 1/20
------
20x + 20*36 = 36x
------------
16x = 20*36
-----
4x = 5*36
x = 5*9
---
x = 45 minutes (time for Otto to do the job alone)
======================
Cheers,
Stan H.